Session: Nov 1999 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Tue Nov 30 10:39:20 MET 1999) Success
No known problems. Weather is rather foggy, so system temperature somewhat higher.

Effelsberg (A. Kraus at Tue Nov 30 11:52:19 MET 1999) Minor failures
At about UT0230 the focus-drive failed, so that the telescope probably had been defocussed since then. For the repair the observations were suspended from UT0707 - 0750. After that (and before UT0230), everything looks good.

Jodrell (Alastair Gunn at Wed Dec 1 03:41:13 MET 1999) Success
No known problems.

Onsala (John Conway at Mon Dec 13 09:14:35 MET 1999) Success
No know problems, winds were near limit though.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:18:47 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: em033a.antab. Read the comments in the file to get further information.